Hello,
I've been working on this problem for hours. This is my last hope of finding an answer. I hope you are able to solve this.
Solve the system of equations. (It's worth 10 points on my assignment)
x - 2y + 3z = 7
2x + y + z = 4
-3x + 2y - 2z = -10
The closest I have come to solving this system was: x = 0, y = 1, z = 3
When I was double checking the solutions,the top two equations checked out to work, but the last equation of the set of three didn't equal out. I don't know if this means that there is no solution, or if I have been doing it completely wrong.
If I could get an answer ASAP that would be amazing! Thank you in advance!!!

Answer by stanbon(75887)

x - 2y + 3z = 7
2x + y + z = 4
-3x + 2y - 2z = -10
-------
I used my TI-84 to get:
x = 2
y = -1
z = 1
=============
Cheers,
Stan H.
==================